Hancock Inc. retains most of its earnings. The company currently has earnings per share of $11. Hancock expects its earnings to grow at a constant rate of 2 percent per year. Furthermore, the average PE ratio of all other firms in Hancock's industry is 14. Hancock is expected to pay dividends per share of $3.50 during each of the next three years. If investors require a 10 percent rate of return on Hancock stock, a fair price for Hancock stock today is $________.

113.95

111.32

131.49

none of the above
Answer #1

EPS in year 3 =EPS*(1+growth)^3 =11*(1+2%)^3 =11.673288
Price in year 3 =EPS*P/E ratio =11.673288*14 =163.4260
Fair Price of Stock =3.50/1.1+3.50/1.1^2+3.50/1.1^3+163.4260/1.1^3 =131.49

Option c is correct option